The atmosphere contains the highly reactive molecule, OH, which acts to remove selected pollutants. Use the values for \Delta H rxn given below to find the \Delta H rxn for the formation of OH and H from water.
1/2 H2(g) + 1/2 O2(g) \rightarrow OH(g) \DeltaHrxn = 42.1 kJ
H2(g)\rightarrow 2H(g) \Delta Hrxn = 435.9 kJ
H2(g) + 1/2 O2(g)\rightarrow H2O(g) \DeltaHrxn = -241.8 kJ
H2O(g) \rightarrow H(g) + OH(g) \DeltaHrxn = ?
1
Expert's answer
2019-11-25T07:25:25-0500

(I) "\frac{1}{2}H_2(g)+ \frac{1}{2}O_2(g) \rightarrow OH(g)" "\Delta H_{rxn} = 42.1 kJ"


(II) "H_2(g) \rightarrow 2H(g)" "\Delta H_{rxn} = 435.9 kJ"


(III) "H_2(g) + \frac{1}{2}O_2(g) \rightarrow H_2O(g)" "\Delta H_{rxn} = -241.8 kJ"


To get "\Delta H_{rxn}" for the formation of OH and H from water we should add "- (III) + (I)+\frac{(II)}{2}"


-(III) "H_2O(g) \rightarrow H_2(g) + \frac{1}{2}O_2(g)" "\Delta H_{rxn} = 241.8 kJ"


(I) "\frac{1}{2}H_2(g) + \frac{1}{2} O_2(g)\rightarrow OH(g)" "\Delta H_{rxn} = 42.1 kJ"


"\frac{(II)}{2}" "\frac {1}{2}H_2 (g) \rightarrow H(g)" "\Delta H_{rxn} = \frac{435.9}{2} = 217.95 kJ"


After addition:


"H_2O(g) + \frac{1}{2}H_2(g)+\frac{1}{2}O_2(g) + \frac{1}{2}H_2(g) \rightarrow H_2(g)+\frac{1}{2}O_2(g) + OH(g)+ H(g)"


"H_2O(g) \rightarrow OH(g) + H(g)" "\Delta H_{rxn} = 241.8 + 42.1 +217.95 =501.85 kJ"